circular wire loop of radius "a" A carries a total charge Q distributed uniformly over its length.A small length dL of the wire is cut off, find the electric field at the centre due to the remaining wire.

Grade:12

2 Answers

vikas askiitian expert
13 years ago

total charge=Q

charge per unit length=Q/2pia

charge contained by dL lenght of element = dq=QdL/2pia ...........1

now

     net electric field at the center of ring is E= 0 ....

when dL element is cut from the ring then let the field of remaining is ER..

so we can say that total electric field at center = ER + electric field due to dL element    

       (total electric field is vector sum of all components)

0=ER + kdq/a2

  ER= -(KQdL/2pia3)

 

 

 

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Total charge=Q
charge per unit length=Q/2πa
charge contained by dL lenght of element = dq = QdL/2πa ...........1
now,
net electric field at the center of ring is E= 0
when dL element is cut from the ring then let the field of remaining is ER
So we can say that total electric field at center = ER + electric field due to dL element
(total electric field is vector sum of all components)
0 = ER + kdq/a2
ER = – (KQdL/2πa3)
 
Thanks and regards,
Kushagra